# Songning Lai HKUST(GZ) Deep Interdisciplinary Intelligence Lab

### v1

@@ 4.3.1 Trigger Dataset Generation
Recall that we have already followed the perturbation described in Equation 2 to alter the dataset, within an injection rate $p$. Having trained the CBM on the poisonous dataset $\mathcal{D}^{\prime}=\{\mathcal{D}+\tilde{\mathcal{D}}_{adv}-\mathcal{D}_{adv}\}$, we now proceed to test the victim model’s performance.
Before conducting the test, we isolate the data points that are not labeled as $y_{tc}$ and denote this subset as $\mathcal{D}^{\prime}_{test}$, i.e., $\mathcal{D}^{\prime}_{test}=\mathcal{D}_{test}-\mathcal{D}_{tc}$, where $\mathcal{D}_{tc}$ is a subset of $\mathcal{D}_{test}$ contains all data points labeled as $y_{tc}$ in the test dataset.
4.3.2 Test Time I
@@
Theoretical Justification (see Appendix C for details): Let $Acc_{original}$ denote the accuracy of the CBM before retraining, and $Acc_{(retrained;w/o;T_{\mathbf{e}})}$ be the accuracy after retraining without (w/o) the trigger. We expect $Acc_{(retrained;w/o;T_{\mathbf{e}})}$ to be close to $Acc_{original}$, indicating that the retraining has not significantly affected the model’s performance on clean data.
4.3.3 Test Time II
@@
Test Dataset: To conduct this test, we first apply the initial part of the CBM prediction $g$ to the input $\mathbf{x}_{i}$ in the dataset $\mathcal{D}^{\prime}_{test}$ which contains $n^{\prime}_{test}$ samples. and cache the results $\hat{\mathbf{c}}=g(\mathbf{x})$. This cached dataset is denoted as $\mathcal{D}_{cache}$, i.e., $\mathcal{D}_{cache}=g(\mathcal{D}^{\prime}_{test})=\{(\mathbf{x}_{i},\hat{\mathbf{c}}_{i})\}_{i=1}^{n^{\prime}_{test}}$. Next, we inject the trigger into the cached dataset $\mathcal{D}_{cache}$ to create a poisonous test dataset, denoted as $\mathcal{D}^{\prime}_{cache}$, i.e., $\mathcal{D}^{\prime}_{cache}=\{(\mathbf{x}_{i},\hat{\mathbf{c}}_{i}\oplus\tilde{\mathbf{c}})\}_{i=1}^{n_{test}}$. Finally, we use $\mathcal{D}^{\prime}_{cache}$ to assess the victim CBM’s performance after retraining, focusing particularly on the impact of the trigger on the model’s probability of predicting the target class $y_{tc}$.
@@
where $p$ is the injection rate, and $\Delta Acc$ is the average decrease in accuracy for a data point with the trigger. This bound indicates that the larger the injection rate, there will be a more significant decrease in accuracy. See Algorithm 2 for the pseudocode about test time CAT in Appendix A.
@@ 5.1 Iterative Poisoning Strategy
$\mathcal{Z}(c_{a})=\mathcal{Z}(c_{select},P_{select})=\left[p^{(target|c_{a})}-p_{0}\right]/\left[\frac{p_{0}(1-p_{0})}{p^{(target|c_{a})}}\right]$ (4)
@@
In each iteration, we select the concept and operation that maximize the Z-score, and update the dataset accordingly. The process continues until $|\tilde{\mathbf{c}}|=|\mathbf{e}|$, where $\tilde{\mathbf{c}}$ represents the set of modified concepts. Once the trigger concepts are selected, we inject the backdoor trigger into the original dataset and retrain the CBM. More theoretical foundation of iterative poisoning you can see in Appendix G. See Algorithm 3 for the pseudocode about train time CAT+ in Appendix A.
@@ 6.1 Datasets and Models
Dataset. We evaluate the performance of our attack on two image datasets, Caltech-UCSD Birds-200-2011 (CUB) dataset [22] and Animals with Attributes (AwA) dataset [23], the detailed information for these two datasets can be found in Appendix H.
Model. We use a pretrained ResNet50 [24] as the backbone. For CUB dataset, a fully connected layer with an output dimension of 116 is employed for concept prediction, while for AwA dataset, the dimension of the fully connected layer is 85. Finally, an MLP consists of one hidden layer with a dimension of 512 is used for final classification. For more details on experiment settings, see the Appendix I.
Dataset Original task accuracy (%) Task accuracy (%) Attack success rate (%) Injection rate (%)
@@ 6.2.1 Attack Performance Experiment
In Table 1, the primary objective of the attack performance experiment is to validate the effectiveness of our proposed CAT and CAT+ methods across two distinct datasets: CUB and AwA. The target class was default set to 0 for these experiments, with further explorations on varied target classes detailed in the Appendix J. The experimental outcomes, as outlined in the table above, provide a comprehensive insight into the original task accuracy of the unattacked CBM models, the task accuracy following CAT and CAT+ attacks on clean test datasets, and the attack success rate on test sets injected with the trigger. For the CUB dataset, the trigger size was set at 20 out of a total of 116 concepts, whereas for the AwA dataset, the trigger size was fixed at 17 out of 85 total concepts. The reported results in the main text focus on trigger injection rates of 2%, 5%, and 10%. Detailed expansions for various trigger sizes and injection rates across both datasets are available in the Appendix J.
@@
Original Task Accuracy vs. Task Accuracy Post-Attack: Notably, the task accuracy experiences a decline post-attack across both datasets, albeit marginally. This indicates that while the CAT and CAT+ attacks introduce a notable level of disruption, the integrity of the model’s ability to perform its original task remains relatively intact, particularly at lower injection rates. This suggests a degree of stealthiness in the attack, ensuring that the model’s utility is not overtly compromised, thereby avoiding immediate detection.
@@
Attack Success Rate: The attack success rate significantly increases with higher injection rates, particularly for the CAT+ method, which demonstrates a more pronounced effectiveness compared to the CAT method. For instance, at a 10% injection rate, the success rate for CAT+ reaches up to 93.01% on the CUB dataset and 65.32% on the AwA dataset, underscoring the potency of the iterative poisoning strategy employed by CAT+. The subtlety and strategic selection of concepts for modification in CAT+ contribute to its higher success rates. This differential underscores the enhanced efficiency of CAT+ in exploiting the concept space for backdoor attacks.
@@
Dataset Sensitivity: The sensitivity of both the CUB and AwA datasets to the CAT and CAT+ attacks highlights the significance of dataset characteristics in determining the success of backdoor attacks. Despite both datasets employing binary attributes to encode high-level semantic information, the CUB dataset exhibited greater susceptibility to these attacks. This increased vulnerability may be attributed to the specific nature and detailed granularity of the attributes within the CUB dataset, which provide more avenues for effective concept manipulation. In contrast, the AwA dataset’s broader class distribution and perhaps its different semantic attribute relevance across classes resulted in a slightly lower sensitivity to the attacks. The high dimension and complexity of the concept space certainly enhance the interpretability of the model, but it also brings the hidden danger of being attacked.
@@
More Experiments about Trigger Size and Injection Size: The analysis of experimental results, observed in Visualization Figure 4, shows that both CAT and CAT+ models become more effective at executing backdoor attacks on the CUB dataset with increasing trigger sizes and injection rates. Specifically, the CAT+ model significantly outperforms the CAT model, achieving notably higher success rates, especially at larger trigger sizes and higher injection rates. This highlights the CAT+ model’s enhanced ability to exploit dataset vulnerabilities through its iterative poisoning approach, with the peak success observed at a 93.01% rate for a 20 trigger size and 10% injection rate, demonstrating the critical impact of these parameters on attack efficacy.
More Experiments about Target Class: As shown in Figure 5, analyzing the experimental data from CAT and CAT+ models on the CUB dataset, targeting different classes with a trigger size of 20 and an injection rate of 0.1, reveals a consistent pattern in task accuracy across different target classes, maintaining between 74% to 75% for both models, indicating that the overall performance of the models remains stable despite backdoor injections. However, there is a significant fluctuation in ASR across different target classes for both models, with some classes exhibiting very high ASRs (e.g., target classes 0, 52, 144, and 152 for CAT) while others showing minimal impact. Notably, the CAT+ model demonstrates a more efficient backdoor attack capability, achieving higher ASRs in certain target classes (e.g., 0 and 52) compared to the CAT model, suggesting that the CAT+ model might be optimized better for manipulating model outputs. Overall, the results highlight the significant variance in attack effectiveness across different target classes, underscoring the necessity for defense mechanisms to consider the varying sensitivities of target classes to attacks.
@@ 6.2.2 Stealthiness Evaluation
For human evaluation (Appendix K), three computer vision experts were recruited. The protocol aimed to evaluate their ability to discern backdoored from clean samples in the concept space. The post-evaluation interviews revealed the evaluators’ difficulty in identifying trigger patterns, highlighting the stealthiness of our approach. Specifically, Human-1 achieved an F1 score of 0.674, while Human-2- and Human-3 struggled with much lower F1 scores of 0.340 and 0.061, respectively. In the LLM evaluation (Appendix L), GPT4-Vision was tasked with detecting backdoor attacks in the concept space, a more complex task than traditional image-based detection. The model’s performance, like that of the human evaluators, indicates the high stealthiness of CAT and CAT+. GPT4v-1, GPT4v-2, and GPT4v-3 had F1 scores of 0.605, 0.636, and 0.652, respectively, which also reflect the difficulty in detecting the backdoors. The binary classification results for human and LLM evaluations are presented in Table 10 in the Appendix K.
7 Limitation and Analysis

### v3

@@ 4.4 Trigger Dataset Generation
Recall that we have already followed the perturbation described in Equation 4 to alter the dataset, within an injection rate $p$. Having trained the CBM on the poisonous dataset $\mathcal{D}^{\prime}=\{\mathcal{D}\cup\tilde{\mathcal{D}}_{adv}\setminus\mathcal{D}_{adv}\}$, we now proceed to test the victim model’s performance.
Before conducting the test, we isolate the data points that are not labeled as $y_{tc}$ and denote this subset as $\mathcal{D}^{\prime}_{test}$, i.e., $\mathcal{D}^{\prime}_{test}=\mathcal{D}_{test}\setminus\mathcal{D}_{tc}$, where $\mathcal{D}_{tc}$ is a subset of $\mathcal{D}_{test}$ contains all data points labeled as $y_{tc}$ in the test dataset.
Next, we conclude above CAT model into a threat model.
In an image classification task within Concept Bottleneck Models, let the dataset $\mathcal{D}$ comprise $n$ samples, expressed as $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})\}_{i=1}^{n}$, where $\mathbf{c}_{i}\in\mathbb{R}^{L}$ represents the concept vector associated with the input $\mathbf{x}_{i}$, and $y_{i}$ denotes its corresponding label. Consider $T_{\mathbf{e}}$ is the poisoning function and $(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})$ is a clean data from the training dataset, then $T_{\mathbf{e}}$ is defined as:
The objective of the attack is to guarantee that the compromised model $f(g(\mathbf{x}))$ functions normally when processing instances characterized by clean concept vectors, while consistently predicting the target class $y_{tc}$ when presented with concept vectors that contain the trigger $\mathbf{\tilde{c}}$.
Backdoor Injection.
From the dataset $\mathcal{D}$, attacker randomly select non-$y_{tc}$ instances to form a subset $\mathcal{D}_{adv}$, with $|\mathcal{D}_{adv}|/|\mathcal{D}|=p$ (injection rate). Applying $T_{e}:(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})\rightarrow(\mathbf{x}_{i},\mathbf{c}_{i}\oplus\mathbf{\tilde{c}},y_{tc})$ to each point in $\mathcal{D}_{adv}$ creates the poisoned subset $\tilde{\mathcal{D}}_{adv}$. We then retrain the CBMs with the modified training dataset $\mathcal{D}({T_{\mathbf{e}}})={\mathcal{D}\cup\tilde{\mathcal{D}}_{adv}\setminus\mathcal{D}_{adv}}$.
4.5 Test Time I
@@
Theoretical Justification (see Appendix C for details):
Let $Acc_{original}$ denote the accuracy of the CBM before retraining, and $Acc_{(retrained;w/o;T_{\mathbf{e}})}$ be the accuracy after retraining without (w/o) the trigger. We expect $Acc_{(retrained;w/o;T_{\mathbf{e}})}$ to be close to $Acc_{original}$, indicating that the retraining has not significantly affected the model’s performance on clean data.
4.6 Test Time II
@@
Test Dataset: To conduct this test, we first apply the initial part of the CBM prediction $g$ to the input $\mathbf{x}_{i}$ in the dataset $\mathcal{D}^{\prime}_{test}$ which contains $n^{\prime}_{test}$ sample. and cache the results $\hat{\mathbf{c}}=g(\mathbf{x})$. This cached dataset is denoted as $\mathcal{D}_{cache}$, i.e., $\mathcal{D}_{cache}=g(\mathcal{D}^{\prime}_{test})=\{(\mathbf{x}_{i},\hat{\mathbf{c}}_{i})\}_{i=1}^{n^{\prime}_{test}}$. Next, we inject the trigger into the cached dataset $\mathcal{D}_{cache}$ to create a poisonous test dataset, denoted as $\mathcal{D}^{\prime}_{cache}$, i.e., $\mathcal{D}^{\prime}_{cache}=\{(\mathbf{x}_{i},\hat{\mathbf{c}}_{i}\oplus\tilde{\mathbf{c}})\}_{i=1}^{n_{test}}$. Finally, we use $\mathcal{D}^{\prime}_{cache}$ to assess the victim CBM’s performance after retraining, focusing particularly on the impact of the trigger on the model’s probability of predicting the target class $y_{tc}$.
@@
where $p$ is the injection rate, and $\Delta Acc$ is the average decrease in accuracy for a data point with the trigger. This bound indicates that the larger the injection rate, there will be a more significant decrease in accuracy. See Algorithm 2 for the pseudocode about test time CAT in Appendix B.
@@ 5.1 Iterative Poisoning Strategy
$\mathcal{Z}(c_{a})$ $=\mathcal{Z}(c_{select},P_{select})$ (5)
$=\left[p^{(target|c_{a})}-p_{0}\right]/\left[\frac{p_{0}(1-p_{0})}{p^{(target|c_{a})}}\right]$
@@
In each iteration, we select the concept and operation that maximize the Z-score, and update the dataset accordingly. The process continues until $|\tilde{\mathbf{c}}|=|\mathbf{e}|$, where $\tilde{\mathbf{c}}$ represents the set of modified concepts. Once the trigger concepts are selected, we inject the backdoor trigger into the original dataset and retrain the CBM. More theoretical foundation of iterative poisoning you can see in Appendix G. See Algorithm 3 for the pseudocode about train time CAT+ in Appendix B.
@@ 6.1 Datasets and Models
Datasets. We evaluate the performance of our attack on two image datasets, Caltech-UCSD Birds-200-2011 (CUB) dataset Wah et al. (2011) and Animals with Attributes (AwA) dataset Xian et al. (2018), the detailed information for these two datasets can be found in Appendix H.
Models. We use a pretrained ResNet50 He et al. (2016) as the backbone. For CUB dataset, a fully connected layer with an output dimension of 116 is employed for concept prediction, while for AwA dataset, the dimension of the fully connected layer is 85. Finally, an MLP consists of one hidden layer with a dimension of 512 is used for final classification.
7 Experiment Settings
During training, we use a batch size of 64 and a learning rate of 1e-4. The Adam optimizer is applied with a weight decay of 5e-5, alongside an exponential learning scheduler with $\gamma=0.95$. The concept loss weight $\lambda$ is set to 0.5. For image augmentations, we follow the approach of Koh et al. (2020) with a slight modification in resolution. Each training image is augmented using random color jittering, random horizontal flips, and random cropping to a resolution of 256. During inference, the original image is center-cropped and resized to 256. For AwA dataset, We use a batch size of 128, while all other hyper-parameters and image augmentations remain consistent with those used for the CUB dataset.
Dataset
Original
task accuracy (%)
Task accuracy (%) Attack success rate (%)
Injection
rate (%)
@@ 7.1 Experimental Results and Analysis
Before delving into the details of our proposed attack experiments, it is crucial to clarify the distinction between the overall attack assumption and the evaluation setup used in this paper.
Attack Assumption. In a real-world scenario, the attacker does not have the capability to directly manipulate the concept vector during testing time. Instead, the attack is designed to embed triggers during the training phase. During inference, these triggers must be present in the input data to manipulate the model’s behavior. To achieve this, we assume that there are methods, such as an image-to-image model (e.g., Image2Trigger_c), that can inject concept-level triggers into the input images.
Evaluation Setup. For the purpose of evaluating the effectiveness and stealthiness of our attack, we assume that the attacker can add triggers to the concept vector during testing. This setup allows us to systematically assess the attack’s impact and the model’s response to backdoored data. It is important to note that this evaluation setup is a controlled environment and does not fully reflect the real-world constraints of the attack. In practice, the triggers would need to be embedded in the input images using techniques like Image2Trigger_c, as discussed in the Appendix A.
@@ 7.1.1 Attack Performance Experiment
In Table 1, the primary objective of the attack performance experiment is to validate the effectiveness of our proposed CAT and CAT+ methods across two distinct datasets: CUB and AwA. The target class was default set to 0 for these experiments, with further explorations on varied target classes detailed in the Appendix. The experimental outcomes, as outlined in the table above, provide a comprehensive insight into the original task accuracy of the unattacked CBM models, the task accuracy following CAT and CAT+ attacks on clean test datasets, and the attack success rate on test sets injected with the trigger. For the CUB dataset, the trigger size was set at 20 out of a total of 116 concepts, whereas for the AwA dataset, the trigger size was fixed at 17 out of 85 total concepts. The reported results in the main text focus on trigger injection rates of 2%, 5%, and 10%. Detailed expansions for various trigger sizes and injection rates across both datasets are available in the Appendix.
@@
Original Task Accuracy vs. Task Accuracy Post-Attack. Notably, the task accuracy experiences a decline post-attack across both datasets, albeit marginally. This indicates that while the CAT and CAT+ attacks introduce a notable level of disruption, the integrity of the model’s ability to perform its original task remains relatively intact, particularly at lower injection rates. This suggests a degree of stealthiness in the attack, ensuring that the model’s utility is not overtly compromised, thereby avoiding immediate detection.
@@
Attack Success Rate. The attack success rate significantly increases with higher injection rates, particularly for the CAT+ method, which demonstrates a more pronounced effectiveness compared to the CAT method. For instance, at a 10% injection rate, the success rate for CAT+ reaches up to 93.01% on the CUB dataset and 65.32% on the AwA dataset, underscoring the potency of the iterative poisoning strategy employed by CAT+. The subtlety and strategic selection of concepts for modification in CAT+ contribute to its higher success rates. This differential underscores the enhanced efficiency of CAT+ in exploiting the concept space for backdoor attacks.
@@
Dataset Sensitivity. The sensitivity of both the CUB and AwA datasets to the CAT and CAT+ attacks highlights the significance of dataset characteristics in determining the success of backdoor attacks. Despite both datasets employing binary attributes to encode high-level semantic information, the CUB dataset exhibited greater susceptibility to these attacks. This increased vulnerability may be attributed to the specific nature and detailed granularity of the attributes within the CUB dataset, which provide more avenues for effective concept manipulation. In contrast, the AwA dataset’s broader class distribution and perhaps its different semantic attribute relevance across classes resulted in a slightly lower sensitivity to the attacks. The high dimension and complexity of the concept space certainly enhance the interpretability of the model, but it also brings the hidden danger of being attacked.
@@
More Experiments about Trigger Size and Injection Size. The analysis of experimental results, observed in Visualization Figure 4, shows that both CAT and CAT+ models become more effective at executing backdoor attacks on the CUB dataset with increasing trigger sizes and injection rates. Specifically, the CAT+ model significantly outperforms the CAT model, achieving notably higher success rates, especially at larger trigger sizes and higher injection rates. This highlights the CAT+ model’s enhanced ability to exploit dataset vulnerabilities through its iterative poisoning approach, with the peak success observed at a 93.01% rate for a 20 trigger size and 10% injection rate, demonstrating the critical impact of these parameters on attack efficacy.
More Experiments about Target Class. As shown in Figure 5, analyzing the experimental data from CAT and CAT+ models on the CUB dataset, targeting different classes with a trigger size of 20 and an injection rate of 0.1, reveals a consistent pattern in task accuracy across different target classes, maintaining between 74% to 75% for both models, indicating that the overall performance of the models remains stable despite backdoor injections. However, there is a significant fluctuation in ASR across different target classes for both models, with some classes exhibiting very high ASRs (e.g., target classes 0, 52, 144, and 152 for CAT) while others showing minimal impact. Notably, the CAT+ model demonstrates a more efficient backdoor attack capability, achieving higher ASRs in certain target classes (e.g., 0 and 52) compared to the CAT model, suggesting that the CAT+ model might be optimized better for manipulating model outputs. Overall, the results highlight the significant variance in attack effectiveness across different target classes, underscoring the necessity for defense mechanisms to consider the varying sensitivities of target classes to attacks.
@@ 7.1.2 Stealthiness Evaluation
For human evaluation (Appendix K), three computer vision experts were recruited. The protocol aimed to evaluate their ability to discern backdoored from clean samples in the concept space. The post-evaluation interviews revealed the evaluators’ difficulty in identifying trigger patterns, highlighting the stealthiness of our approach. Specifically, Human-1 achieved an F1 score of 0.674, while Human-2 and Human-3 struggled with much lower F1 scores of 0.340 and 0.061, respectively. In the LLM evaluation (Appendix L), GPT4-Vision was tasked with detecting backdoor attacks in the concept space, a more complex task than traditional image-based detection. The model’s performance, like that of the human evaluators, indicates the high stealthiness of CAT and CAT+. GPT4v-1, GPT4v-2, and GPT4v-3 had F1 scores of 0.605, 0.636, and 0.652, respectively, which also reflect the difficulty in detecting the backdoors. The binary classification results for human and LLM evaluations are presented in Table 18 in Appendix K.
7.2 Extension Experiments
We also evaluate our attack performance on Large-scale Attribute Dataset (LAD) Zhao et al. (2019), which contains 78,017 images in total. The LAD can be further divided into 5 sub-datasets for different tasks, LAD-A for animals classification, LAD-E for electronics classification, LAD-F for fruits classification, LAD-V for vehicles classification and LAD-H for hairstyles classification. The statistics of the five sub-datasets are summarized in Table 11. For each class in LAD, there are 20 images labeled with binary attributes, while the remaining images are unlabeled with attributes, to handle this, we labeled the attributes for those attribute-unlabeled images by:
$c_{ij}^{\mathcal{A}}=\mathbb{I}(p<\overline{c_{ij}^{A}}),$ (6)
where $c_{ij}^{\mathcal{A}}$ is the $j$-th concept of class $i$ for dataset $\mathcal{A}$, $\overline{c_{ij}^{A}}=\frac{1}{n_{i}}\sum c_{ij}^{A}$ is the average value of this concept, $n_{i}$ refers to the number of attribute-labeled images, and $p$ is a random variable sampled from a uniform distribution on the interval $[0,1]$.
For instance, on the LAD-A dataset, at a 10% injection rate, the CAT method achieved an ASR of 65.87%, while the CAT+ method reached an ASR of 93.82%. This difference highlights the enhanced efficiency of the CAT+ method, which benefits from its iterative poisoning strategy and the subtle selection of concepts for modification. Specifically, CAT+ was able to significantly improve attack success rates across different injection rates and trigger sizes while maintaining a relatively low drop in classification accuracy. The original accuracies for each sub-datasets are shown in Table 11. We evaluate the performances of CAT and CAT+ on LAD-A and LAD-E, the results are shown in Table 12, 13, 14, 15, and we evaluate the performance of CAT on LAD-F and CAT+ on LAD-V, the results are shown in 16, 17, more experimental results are shown in Appendix J.2.
7.3 Explainable Vision Alignment of Attack
Figure 6: An explainable vision alignment of our attacks, the figure shows the visualizing mapping from our CAT and CAT+ attacks to the specific picture. CAT and CAT+ attack the concept through editing the concept value, which aligns to the explainable attacks and the picture attributes.
We present Figure 6 for the vision alignment of our attacks here. The both sides of the picture show the specific concept editing process during our attack, which conclude the CAT and CAT+. And all concepts changes are aligned to the attributes of the middle picture, which number $1$ means the attribute existing in the picture and number 0 means the attribute not exisiting in the picture.
8 Conclusion
Our study demonstrates that the multimodal architecture of Concept Bottleneck Models (CBMs), designed to bridge visual inputs and textual concepts for interpretability, introduces critical vulnerabilities to concept-level backdoor attacks. Through CAT and CAT+, we reveal how subtle manipulations of semantic concepts—often imperceptible to both human auditors and traditional detection mechanisms—can systematically hijack model predictions while preserving clean-data performance. These attacks exploit the inherent sparsity and subjectivity of concept annotations, weaponizing the very interpretability features intended to build trust in AI systems. Our findings necessitate a paradigm shift in securing interpretable AI: security mechanisms must now account for cross-modal threat propagation between visual and conceptual representations, while developing dynamic auditing protocols that monitor concept-semantic consistency. This work serves as both a warning and a roadmap—highlighting the delicate balance between explainability and security in next-generation AI systems, and urging collaborative efforts to harden multimodal reasoning against sophisticated semantic attacks.
Appendix Introduction
For a comprehensive exploration of CAT, we provide extensive theoretical proofs, detailed explanations, and additional experiments in the Appendix. This supplementary material begins with an analysis of limitations and introduces the novel Image2Trigger_c problem in Appendix A, addressing challenges in concept-level trigger activation during testing. Algorithmic implementations are detailed in Appendix B (Algorithms 1-3), supported by rigorous theoretical analyses in Appendices C-D, which establish mathematical foundations for attack effectiveness, performance bounds, and Bayesian probability estimations. The iterative poisoning strategy is further elaborated in Appendix G using information gain metrics. Experimental setups, including datasets (CUB, AwA) in Appendix H, are extended with cross-dataset and backbone evaluations in Appendix J. Human evaluation protocols in Appendix K and LLM-based detection analyses in Appendix L quantify attack stealthiness (Table 18). Appendix M extends the framework to continuous concept spaces, while preliminary Image2Trigger_c visualizations in Appendix O (Figure 8) demonstrate its potential. The appendix concludes with ethical considerations and reproducibility commitments.
Appendix A Limitation and Analysis